OpenStudy (anonymous):

if y varies inversely as the square of x and y=50 when x=4, find y when x=5

OpenStudy (jziggy):

y varying inversely to the square of x is saying that y = k/x^2.

OpenStudy (jziggy):

So, plug in your first set of numbers and find k

OpenStudy (anonymous):

ok I got that part. my answer was 160

OpenStudy (acxbox22):

50=k/(4^2) 50=k/16 k=800<--- this was the first step now y=800/(5^2) y=800/25 y=??

OpenStudy (anonymous):

huh, well y would be 32 but that's not the answer I got

OpenStudy (acxbox22):

you forgot to square the 5 and just divided 800 by 5 instead of 25

OpenStudy (anonymous):

can you find my mistake? y(x^2) =y(x^2) 50(4^2)/5=y(5^2)/5 y=160

OpenStudy (jziggy):

tbh, not even sure what you're trying to do with that :/ \[y =(k/x^2) --> 50 = k/16 --> k = 800\] \[y = k/x^2 ---> y = 800/5^2 ---> y = 800/25\]

OpenStudy (anonymous):

ok, I guess I understand

OpenStudy (anonymous):

will you help me with 3 more